OpenStudy (anonymous):

If sin(theta) = -2/3, which of the following are possible?

OpenStudy (anonymous):

i guess lots of things are possible are you looking for the cosine?

OpenStudy (anonymous):

OH! Gomenasai! I forgot to specify...

OpenStudy (anonymous):

lol i had to look that up

OpenStudy (anonymous):

my japanese is limited to "saki" "samurai" and "unami"

OpenStudy (anonymous):

A. \[\sec \theta = -3/2 and \tan \theta = 2/\sqrt{5}\]

OpenStudy (anonymous):

slow

OpenStudy (anonymous):

don't post any more answers

OpenStudy (anonymous):

lets draw a triangle with one side 2 and the hypotenuse 3 like this 

OpenStudy (anonymous):

there is a picture of an angle whose sine is \(\frac{2}{3}\) as in "opposite over hypotenuse" don't worry about it being negative yet

OpenStudy (anonymous):

what you need is the third side, which you find via pythgoras

OpenStudy (anonymous):

Ah, the Pythagorean theorem.

OpenStudy (anonymous):

let me know when you get \(a=\sqrt5\)

OpenStudy (anonymous):

I'm terrible at this, and I apologize for that.

OpenStudy (anonymous):

\[a^2+2^2=3^2\] solve for \(a\)

OpenStudy (anonymous):

you can do it in your head what is 3 squared?

OpenStudy (anonymous):

9.

OpenStudy (anonymous):

what is 2 squared?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

And 1 squared, being as it's 1 * itself would be 1, yes?

OpenStudy (anonymous):

Wait a minute...

OpenStudy (anonymous):

what is 9 - 4?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

so \[a=\sqrt{3^2-2^2}=\sqrt{9-4}=\sqrt5\] that is all

OpenStudy (anonymous):

So A^2=5, ergo A=2.5

OpenStudy (anonymous):



OpenStudy (anonymous):

ergo my foot the square root of 5 is not half of 5!

OpenStudy (anonymous):

I get it, because the Pythagorean theorem was square values. We have to fin the regular values. Sorry, demo it is getting really late and I can't think right.